# A Bag of Tricks for Few-Shot Class-Incremental Learning

### v1

@@ 3.4 Adaptability Tricks
While stability tricks help the model retain the knowledge of base classes, they offer limited adaptability for the model to learn novel classes effectively. Independent performance evaluation of base and novel classes at the end of training demonstrates that the performance of novel classes is substantially inferior to that of base classes [33]. Consequently, the overall performance at the end of training is predominantly influenced by the performance on the base classes. This underscores the need for FSCIL methods to enhance the model’s adaptability on incremental sessions to improve the model’s performance on novel classes. In this section, we discuss two tricks that provide more adaptability for the model: incremental fine-tuning and SubNet tuning.
@@
Incremental Fine-tuning. Fine-tuning is a common practice in machine learning literature to tune a pre-trained model for a new task or setting. It is also a widely used technique across conventional continual learning literature, predominantly with rehearsal-based techniques. However, fine-tuning in the context of FSCIL requires careful consideration of the training setup, since no data from the previous session is available for rehearsal during the current session. Consequently, tuning can cause catastrophic forgetting of already learned knowledge. In our framework, we adopt the fine-tuning concept of [33], which utilizes a small learning rate to tune certain portions of the pre-trained encoder, while keeping the rest of the encoder frozen. Specifically, we freeze the shallow layers of the network since shallower layers are known to capture domain-invariant features, whereas the deeper layers learn more fine-grained features.
@@
SubNet Tuning While the incremental fine-tuning approach (previous trick) provides more adaptability for learning novel classes, it may result in decreased stability. To deal with this issue, we introduce our next adaptability trick called SubNet tuning, which maintains the adaptability of incremental fine-tuning without a substantial drop in stability. The concept is inspired by the Lottery-ticket hypothesis [10], which states that there exists a subnetwork of a dense network that performs as well as the whole network. In [14], this concept is utilized in the context of FSCIL to find a sub-network of the trained encoder that performs on par with the whole network on the base classes. Adopting this into our framework, we freeze the sub-network during incremental training and allow the remaining parameters to be tuned for learning the novel classes. Following [14], we extract a SubNet (subnetwork) mask ${m}^{*}$ as:
$\small{m}^{*}=\underset{{m}\in[0,1]^{|{\theta}|}}{min}~{}\frac{1}{n}\sum^{n}_{i=1}\mathcal{L}\big{(}f_{({\theta}\odot{m})}({x}_{i}),y_{i}\big{)}-\mathcal{L}\big{(}f_{\theta}({x}_{i}),y_{i}\big{)}.$ (2)
@@
Here, $\mathcal{L}$ is the loss function for training the model on the base session, and $m^{*}$ is the optimal binary mask with the same size as the network, for which the performance of the masked SubNet is comparable to that of the original network.
@@ 3.5 Training Tricks
Building upon the principles of stability and adaptability tricks discussed earlier, we introduce a set of training techniques that can further enhance overall performance without compromising either stability or flexibility. These tricks include training a larger encoder, adding a pre-training step, and including an additional learning signal.
Larger Encoder. Most existing works on FSCIL use relatively shallow encoders like ResNet-18 and ResNet-20 as the backbone [33, 34]. Since the amount of training data is exceedingly scarce in FSCIL, even for the base session, larger encoders face overfitting and loss of stability. In our framework, we explicitly deal with the stability problem by incorporating the three aforementioned stability tricks (Sections 3.3) that enable us to train larger encoders. As a result, we train an encoder with approximately twice the parameters compared to prior works.
@@
[…]
@@
Including Additional Learning Signal. Following the intuition from the previous trick, we include another self-supervised learning signal, but this time while training on the base session rather than as a separate step. Existing literature on other data-scarce scenarios, such as semi-supervised learning, has shown that adding a pre-text task along with the supervised learning helps the model learn better representation without overfitting to the small labelled set [1, 28]. To this end, we include a rotation prediction task [11] that is shown to perform well with supervised learning [1]. Here, the basic idea is to apply a rotation operation on the input image, and the task is to predict the amount of rotation applied to the image. In practice, a rotation module randomly samples one of the following rotations and applies it to the image: $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$. As a result, the rotation prediction task can be viewed as a four-way classification task, represented as:
$\small\mathcal{L}_{rot}=\mathcal{H}(r,P_{\theta_{r}}(r|Rotate(x)).$ (4)
@@ 4.1 Datasets and Implementation Details
Following the established protocol in the FSCIL literature, we conduct our experiments on three popular datasets: CIFAR100 [16], miniImageNet [30] and CUB200 [35]. To ensure a fair comparison with prior works on FSCIL, we use the same encoder (ResNet-18), and data split across the training sessions as in [34, 8]. Specifically, for CIFAR-100 and miniImageNet, we use 60 classes for the base session and 40 classes for the incremental sessions. The incremental learning experiments are conducted on a 5-way, 5-shot setting. In the case of CUB-200, we allocate 100 classes for the base session and another 100 classes for the incremental sessions, each containing ten classes (10-way, 5-shot). Further details on implementation and hyper-parameters are presented in Appendix S1.
4.2 Main Results
Baselines. As outlined in Section 3.2, we adopt the incremental-frozen framework as the baseline for our study. As shown in Table 1, the accuracy of this
Table 1: Impact of different tricks on the baseline for all datasets.
Setting CIFAR-100 CUB-200 miniIN
Baseline 43.77 59.88 45.08
+ SupCon 50.16 60.38 48.90
+ ETF vector 51.10 60.74 49.73
+ Pseudo-classes 51.65 62.18 54.79
+ Finetuning 52.51 62.27 55.82
+ SubNet tuning 58.12 63.10 57.85
+ Larger encoder 58.52 63.55 57.99
+ Pre-training 58.62 63.59 58.10
+ Additional signal 58.75 63.75 59.57
baseline for CIFAR-100, CUB-200, and miniImageNet datasets are 43.77%, 59.88%, and 45.08%, respectively. To ensure the best performance for the baseline, we perform an extensive hyper-parameter study presented in Appendix S4.1. As we later compare our results with prior works (Section 4.4), we observe that this baseline outperforms many of the prior studies, showing the robustness of our baseline.
Stability Tricks. Next, we discuss the results of including stability tricks into the baseline, beginning with the addition of SupCon loss. The results from this study show substantial improvement for all datasets (Table 1), obtaining accuracies of 50.16%, 60.38%, and 48.90%, for CIFAR-100, CUB-200, and miniImageNet respectively. As discussed earlier, a key element for ensuring high stability in FSCIL is to ensure increased separability in the embedding space (increased inter-class distance and reduced intra-class dispersion). In Figures 1(a) and 1(b), we plot the cumulative probability of inter-class and intra-class distances (defined in Appendix S2 for different tricks. As we observe from Figure 1(a), adding SupCon provides a large increase in the inter-class distance compared to training with cross-entropy loss in the baseline. At the same time, SupCon greatly reduces the intra-class distance compared to the cross-entropy-based baseline (illustrated in Figure 1(b). However, our findings diverge from those of [33], which indicated that while supervised contrastive loss effectively reduces intra-class distances, it unexpectedly leads to a reduction in inter-class distances as well. The divergence in our findings may be attributed to implementation specifics. For example, SAVC utilized multi-crop augmentation, which we do not use. Also, we adopted a different set of hyper-parameters determined through our study (Appendix S4.2).
We also show the overall class separation degree for different tricks in Figure 1(c). Class separation degree refers to the degree of distinctiveness or separability between different categories within the embedding space, which can be measured as defined in Appendix S2. It typically ranges between 0 and 1, with higher values indicating clearer boundaries and better separability between classes. As we observe, adding SupCon provides better class separation for both base and novel classes. Overall, the higher separation in the embedding space contributes to the large improvement in the overall accuracy. Analyzing the performance of both base and novel classes in Figure 1(d), we find improvements for both, with a 5.4% boost in base class accuracy and a 7.8% increase in novel class accuracy.
Our first adaptability trick, incremental fine-tuning, provides adaptability by means of fine-tuning the encoder in the incremental sessions. As shown in Figure 2(a), this trick improves accuracy on novel classes by 9.2%. We also show the accuracy after each session in Figure 2(b). From this figure, we observe comparable accuracies to that of stability tricks in the earlier sessions. However, fine-tuning shows better performance in later sessions when the number of novel classes increases. This again shows the importance of the fine-tuning trick for learning novel classes. Overall, this trick provides an improved accuracy of 52.51%, 62.37% and 55.82% on the CIFAR-100, CUB-200, and miniImageNet, respectively.
While the fine-tuning trick improves the overall performance, it causes the model to forget some of its already-learned knowledge, which is evident from
the performance of base classes in Figure 2(a). Here, the performance on the base classes drops by 4.7% when we include the fine-tuning trick. We resolve this issue with our next adaptability trick, SubNet tuning. As we see from Figure 2(a), SubNet tuning achieves comparable accuracy to that of the fine-tuning trick on novel classes. However, unlike the fine-tuning trick, the improvements on novel classes do not come at the cost of a drop in base classes. In fact, we get an improved accuracy of 76.19% on the novel classes, and 58.1% on total classes.
Analyzing the performance over the incremental sessions in Figure 2(b), we observe that the accuracy for SubNet tuning is constantly higher than the previous tricks by a large margin. In terms of overall performance, we find an improved accuracy of 58.12%, 63.1%, and 57.85% for the CIFAR-100, CUB-200, and miniImageNet datasets. In Figure 2(c), we visualize a t-SNE representation of the learned embeddings (for a randomly selected subset of classes) for the stability tricks and the adaptability tricks. As observed in these figures, stability tricks alone provide sufficient space for the integration of novel classes (5 and 6), although it results in a slight overlap with base classes. In contrast, the incorporation of adaptability tricks creates a more separable embedding space.
(a) Baseline.
(b) bag of tricks.
Training Tricks. Finally, we discuss the results of adding the training tricks. First, we investigate the impact of training a larger encoder by increasing the encoder size from 11.2M (ResNet-18) to 25.6M (ResNet-50). With the larger encoder, the overall accuracy increases to 58.52%, 63.55% and 57.99%, respectively, on the three datasets. To ensure fairness in comparison to prior works, we investigate the performance of prior state-of-the-art methods, SAVC [33] and SoftNet [14], with the larger encoder (see Appendix S4.4). This experiment reveals a substantial decline in performance across all datasets, underscoring the inability to effectively train a larger encoder without the stability tricks.
Next, adding the pre-training trick yields additional improvements. With this trick, the performance on the three datasets is increased to 58.52%, 63.55%, and 57.99%, respectively. Finally, we observe relatively large improvements with our last training trick of incorporating the additional learning signal, resulting in accuracies of 58.75%, 63.75%, and 59.57%, on CIFAR-100, CUB-200, and miniImageNet, respectively. To provide an overall evaluation, we present the confusion matrices for both the baseline and the complete bag of tricks in Figure 4. For the baseline, noticeable dark points (correct predictions) are observed along the diagonal for the base classes (first 60 classes), but these become less prominent for the incremental classes. With the bag of tricks, we observe significant improvements for all classes, including novel ones.
4.3 Ablation Study
Table 2: Ablation study.
Stab. Adap. Train. Acc
58.75
57.85
55.56
46.25
54.79
37.34
38.15
43.77
In this section, we present an ablation study on the stability, adaptability and training tricks on the CIFAR-100 dataset. The results of this study are presented in Table 2. It is evident from the table that stability tricks have the highest individual impact on the performance of the model, removing which results in a 12.5% drop in performance. Furthermore, removing stability tricks with any of the other tricks drastically drops the performance, resulting in a final accuracy worse than the baseline. This is caused by the fact that without stability tricks, adaptability tricks alone overfit the new classes, and training tricks alone (with a larger encoder) do not learn generalized features. We find the second important component to be the adaptability tricks, which results in a 3.19% drop in performance when removed. Finally, removing training tricks shows a 0.9% drop in performance.
Method Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8
Baseline 70.05 65.37 60.59 56.72 53.38 50.53 48.07 45.87 43.77
Rebalancing [13] 61.31 47.80 39.31 31.91 25.68 21.35 18.67 17.24 14.17
iCaRL [26] 61.31 46.32 42.94 37.63 30.49 24.00 20.89 18.80 17.21
TOPIC [34] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42
IDLVQ-C [3] 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84
FSLL [22] 66.48 61.75 58.16 54.16 51.10 48.53 46.54 44.20 42.28
FSLL+SS [22] 68.85 63.14 59.24 55.23 52.24 49.65 47.74 45.23 43.92
F2M [31] 67.28 63.80 60.38 57.06 54.08 51.39 48.82 46.58 44.65
CEC [37] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14
MetaFSCIL [8] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97
CLOM [44] 74.20 69.83 66.17 62.39 59.26 56.48 54.36 52.16 50.25
C-FSCIL [12] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47
LIMIT [41] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23
FACT [40] 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10
SAVC [33] 79.85 73.70 69.37 65.28 61.91 59.27 57.24 54.97 53.12
ALICE [25] 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10
SoftNet [14] 79.88 75.54 71.64 67.47 64.45 61.09 59.07 57.29 55.33
Ours 80.25 77.20 75.09 70.82 67.83 64.86 62.73 60.52 58.75
**We do not compare with NC-FSCIL [36] since it deviates (uses a memory bank) from the standard training protocol for FSCIL.
4.4 Comparison to Prior Works
Finally, we compare the results of our bag of tricks with the prior works on CIFAR-100 in Table 3, and present the results for CUB-200 and miniImageNet in Appendix Tables S1 and S2. The tables present the results for the base session and the overall performance across incremental sessions. The first row in this table presents the results for our baseline. Among prior works, SoftNet held the previous state-of-the-art for CIFAR-100 with a final accuracy of 55.33%. Our framework achieves a performance of 58.75%, marking a 3.42% improvement over the current state-of-the-art. In Figure 5, we plot the accuracy over the incremental sessions for CIFAR-100, CUB-200, and miniImageNet. For CUB-200, the previous state-of-the-art was held by SAVC with an accuracy of 62.50%, which our framework outperforms with an accuracy of 63.75%. Finally, the state-of-the-art on miniImageNet was also held by SAVC, with an accuracy of 57.11%, which our method outperforms by 2.46%, achieving an accuracy of 59.57%.
Performance Across Different Shots. Following prior works in FSCIL [40, 34], we report the main results in a 5-shot setting. Nonetheless, We also investigate the performance on 1-shot and 2-shot settings to evaluate how the model performs in data-scarce settings and compare the performance with two of the previous state-of-the-art methods, SAVC and SoftNet. We also investigate the 10-shot setting to determine how an increase in data impacts performance. As seen from Figure 6, in 1-shot learning, our proposed framework outperforms prior works by a larger margin than the 5-shot setting. This further shows the effectiveness of the bag of tricks in data-scarce settings. More experiments on 1-, 2-, 5- and 10-shot settings are discussed in Appendix S4.5.
Performance on ImageNet-1K. So far we have report the main results of our approach on, CIFAR-100, CUB-200, and miniImageNet, which are commonly used by prior works in this area. However, these datasets are relatively small in terms of the number of classes. To better understand the scalability of FSCIL in terms of the number of classes, we conduct an experiment on the ImageNet-1K [29] dataset, which contains 1000 classes. For this study, we consider 500 randomly selected classes as the base classes and report the results for a 50-way, 10-shot incremental learning setting. The results of this study with our method and its comparison to prior methods are presented in Table 4. As seen from the results, our proposed framework outperforms prior works on the ImageNet-1K dataset by a considerable margin. More specifically, our framework achieves 7.04% and 2.13% improvement over SoftNet and SAVC, respectively.
Method Acc. in each session (%) ↑
0 1 2 3 4 5 6 7 8 9 10
SAVC [33] 37.74 35.94 34.45 33.16 31.88 30.93 29.80 28.83 28.1 27.37 26.12
SoftNet [14] 36.21 34.23 32.35 30.53 28.71 27.15 25.68 24.31 23.16 22.19 21.21
Ours 39.34 37.86 36.49 35.23 34.30 33.83 32.29 31.21 30.77 29.34 28.25
5 Conclusion
We present a bag of tricks framework that combines eight effective tricks in three distinct categories to improve the stability, adaptability, and overall performance of FSCIL. Stability tricks improve the separation among learned classes to facilitate the learning of new ones, resulting in large improvements on both the base and total classes. Adaptability tricks improve the performance on novel classes by providing more learning capability during incremental sessions. Finally, training tricks provide an additional boost to the final performance. Our detailed analysis of each of these tricks shows the effectiveness of each technique, while our framework sets new state-of-the-art on three benchmark datasets.
Limitations. Our framework has a few limitations. Though our framework provides adequate adaptability, the performance of the novel classes is still relatively lower compared to the base classes. Additionally, our proposed framework demands a relatively higher computational budget, primarily due to the training tricks, such as the pre-training step and a larger encoder. Another limitation of our work is that we only included tricks that can be combined with all other tricks into a single framework. Some other tricks, such as the notion of meta-learning, or weight-space manipulation, could not easily be integrated into our framework. Future work can explore certain adaptations or modifications of these techniques that could facilitate their inclusion in the bag of tracks framework.
Supplementary Material
  • S1: Datasets and Implementation Details
  • S2: Evaluation Protocols
  • S3: Comparison to SOTA
  • S4: Additional Results
    • S4.1: Baseline
    • S4.2: Stability Tricks
    • S4.3: Adaptability Tricks
    • S4.4: Training Tricks
    • S4.5: Performance on Different Shots
  • S5: Time Complexity
S1 Datasets and Implementation Details
Following previous works [34, 33], we evaluate our framework on three popular datasets: CIFAR100 [16], miniImageNet [30], and CUB200 [35]. CIFAR-100 is a dataset of 100 classes, where we use 60 classes (following [34]) in the base session and the remaining 40 classes in the incremental sessions. Each incremental session is formulated as a 5-way, 5-shot problem. CUB-200 is a dataset with 200 fine-grained categories. For this dataset, we use 100 classes (following [34]) in the base session and the remaining 100 classes in the incremental sessions, with 10 classes in each session. Finally, miniImageNet is a subset of the popular ImageNet dataset that contains 100 classes. For this dataset, we use 60 classes (following [34]) for the base session and 5 classes per incremental session over 8 sessions.
For the encoder, we use ResNet-18 by default for all datasets, similar to [34, 14]. We train the model with an SGD optimizer, a momentum of 0.9, and a batch size of 64. The learning rate is set to 0.1 for CIFAR-100 and miniImageNet and 0.001 for CUB-200. For all experiments, the model is trained on an Nvidia RTX 2080 GPU.
S2 Evaluation Protocols
Following the standard evaluation protocol in FSCIL literature [34, 33, 14], we report the model’s accuracy after each incremental session. To further understand the properties of the learned representations, we investigate inter-class distance, intra-class distance, and class separation. Following, we define these properties.

### v2

@@ 3.4 Adaptability Tricks
While stability tricks help the model retain the knowledge of base classes, they offer limited adaptability for the model to learn novel classes effectively. Independent performance evaluation of base and novel classes at the end of training demonstrates that the performance of novel classes is substantially inferior to that of base classes (Song et al., 2023). Consequently, the overall performance at the end of training is predominantly influenced by the performance of the base classes. This underscores the need for FSCIL methods to enhance the model’s adaptability on incremental sessions to improve the model’s performance on novel classes. In this section, we discuss two tricks that provide more adaptability for the model: incremental fine-tuning and SubNet tuning, which we combinedly denote as Incremental SubNet Tuning.
@@
Incremental SubNet Tuning. Fine-tuning is a common practice in machine learning literature to tune a pre-trained model for a new task or setting. It is also a widely used technique across conventional continual learning literature, predominantly with rehearsal-based techniques. However, fine-tuning in the context of FSCIL requires careful consideration of the training setup, since no data from the previous session is available for rehearsal during the current session. Consequently, tuning can cause catastrophic forgetting of already learned knowledge. In our framework, we adopt the fine-tuning concept of Song et al. (2023), which utilizes a small learning rate to tune certain portions of the pre-trained encoder, while keeping the rest of the encoder frozen. Specifically, we freeze the shallow layers of the network since shallower layers are known to capture domain-invariant features, whereas the deeper layers learn more fine-grained features.
@@
While such an incremental fine-tuning approach provides more adaptability for learning novel classes, it may result in decreased stability. To deal with this issue, we combine this adaptability trick with the concept of SubNet tuning, which maintains the adaptability of incremental fine-tuning without a substantial drop in stability. The concept is inspired by the Lottery-ticket hypothesis (Frankle & Carbin, 2018), which states that there exists a subnetwork of a dense network that performs as well as the whole network. In Kang et al. (2023), this concept is utilized in the context of FSCIL to find a sub-network of the trained encoder that performs on par with the whole network on the base classes. Adopting this into our framework, after the training of the encoder ($f_{\theta}$) on the base session, we find a subset of network parameter that performs as well as the whole network on the base session.Specifically, we follow Kang et al. (2023) to extract a SubNet (subnetwork) mask ${m}^{*}$ as:
${m}^{*}=\underset{{m}\in[0,1]^{|{\theta}|}}{min}~{}\frac{1}{n}\sum^{n}_{i=1}\mathcal{L}\big{(}f_{({\theta}\odot{m})}({x}_{i}),y_{i}\big{)}-\mathcal{L}\big{(}f_{\theta}({x}_{i}),y_{i}\big{)},$ (3)
@@
where $\odot$ is an element-size dot-product operation. Here, $\mathcal{L}$ is the loss function for training the model on the base session, and $m^{*}$ is the optimal binary mask with the same size as the network, for which the performance of the masked SubNet is comparable to that of the original network. During the incremental session, we freeze the sub-network to ensure the performance of the base session and fine-tune the remaining parameters for learning the novel classes.
@@ 3.5 Training Tricks
Building upon the principles of stability and adaptability tricks discussed earlier, we introduce a set of training techniques that can further enhance overall performance without compromising either stability or adaptability. These tricks include adding a pre-training step and including an additional learning signal.
@@
$\mathcal{L}_{\text{{con}}}=-\frac{1}{2b}\sum_{i=1}^{2b}log\frac{exp(z_{i},z_{\kappa(i)}/\tau)}{\sum_{k=1}^{2b}\mathbf{1}_{[k\neq i]}exp(z_{i},z_{k}/\tau)},$ (4)
@@
Including Additional Learning Signal. Following the intuition from the previous trick, we include another self-supervised learning signal, but this time while training on the base session rather than as a separate step. Existing literature on other data-scarce scenarios, such as semi-supervised learning, has shown that adding a pre-text task along with the supervised learning helps the model learn better representation without overfitting to the small labelled set (Berthelot et al., 2019; Roy & Etemad, 2023). To this end, we include a rotation prediction task (Gidaris et al., 2018) that is shown to perform well with supervised learning (Berthelot et al., 2019). Here, the basic idea is to apply a rotation operation on the input image, and the task is to predict the amount of rotation applied to the image. In practice, a rotation module randomly samples one of the following rotations and applies it to the image: $0^{\circ},90^{\circ},180^{\circ},270^{\circ}$. As a result, the rotation prediction task can be viewed as a four-way classification task, represented as:
$\mathcal{L}_{rot}=\mathcal{H}(r,P_{\theta_{r}}(r|Rotate(x)).$ (5)
@@ 4.1 Datasets and Implementation Details
Following the established protocol in the FSCIL literature, we conduct our experiments on three popular datasets: CIFAR100 (Krizhevsky et al., 2009), miniImageNet (Russakovsky et al., 2015) and CUB200 (Wah et al., 2011). To ensure a fair comparison with prior works (Tao et al., 2020; Chi et al., 2022) on FSCIL, we use the same encoder (ResNet-18), and data split across the training sessions. Specifically, for CIFAR-100 and miniImageNet, we use 60 classes for the base session and 40 classes for the incremental sessions. The incremental learning experiments are conducted on a 5-way, 5-shot setting. In the case of CUB-200, we allocate 100 classes for the base session and another 100 classes for the incremental sessions, each containing ten classes (10-way, 5-shot). Further details on implementation and hyper-parameters are presented in Appendix S1.
4.2 Evaluation Protocols
Following the standard evaluation protocol in FSCIL literature (Tao et al., 2020; Song et al., 2023; Kang et al., 2023), we report the model’s accuracy after each incremental session. To further understand the properties of the learned representations, we investigate inter-class distance, intra-class distance, and class separation. Following, we define these properties.